OpenStudy (anonymous):

Evaluate the integral from 0 to pi/4 for [(1/(cos(x))squared) + sin (2x)] dx

OpenStudy (anonymous):

\[\frac{1}{\cos^2(x)}=\sec^2(x)\] and an anti derivative for secant squared is \[\tan(x)\] is a start

OpenStudy (anonymous):

\[\sin(2x)=2\sin(x)\cos(x)\] and that has \[\sin^2(x)\]as an anti-derivative

OpenStudy (anonymous):

so you have \[\tan(x)+\sin^2(x)\] as your anti derivative. evaluate at upper limit, lower limit and subtract

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/4} [(\frac{1}{\cos(x)^{2}} + \sin (2x)] dx= Tan(x)-\frac{1}{2}Cos(2x)\] Put in the limits of integration and Bingo.....so \[\frac{1}{2}+1=\frac{3}{2}\]

OpenStudy (anonymous):

thanks!